# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v2

@@ A.6 Choosing $\epsilon$ for converting Laplacians to SPD matrices
A.7 Comparison of Two Different Riemannian Metrics for SPD Matrices
The two most commonly used Riemannian metrics on the SPD manifold are the Affine Invariant Riemannian Metric (AIRM) and the Log-Euclidean Riemannian Metric (LERM) (Ilea et al., 2018; Thanwerdas & Pennec, 2023; Chen et al., 2024). AIRM is a Riemannian metric that remains invariant under affine transformations, meaning the metric is unaffected when matrices are transformed by any invertible operation. The geodesic distance between two SPD matrices, A and B, using AIRM is given by (You & Park, 2021; Moakher, 2005):
@@ A.7 Comparison of Two Different Riemannian Metrics for SPD Matrices
On the other hand, LERM addresses some of the computational complexity challenges associated with AIRM by mapping SPD matrices to an Euclidean space through the matrix logarithm operation. In this Euclidean space, computations are simplified. The geodesic distance between two SPD matrices, A and B, using LERM is defined as (Huang et al., 2014):
@@
A.8 Using Normalized Laplacians for GGD Calculation
In many existing studies, the normalized Laplacian matrix is commonly used to study spectral graph properties (Chung, 1997). The normalized Laplacian matrix of a graph $G$ is defined as: $L_{norm}=I-A_{norm}$, where $A_{norm}$ is the normalized adjacency matrix. The normalized adjacency matrix is expressed as (Chung, 1997):
@@ A.8 Using Normalized Laplacians for GGD Calculation
GGD using normalized Laplacian matrices 384.1 254.44 188.345 165.332
A.9 Effect of tuning parameter $\eta$ on Graph Matching
In the original work, it was suggested that the regularization parameter $\eta$ needs to be chosen so that $\sigma\vee n^{-0.1}\lesssim\eta\lesssim 1/\log n$ (Fan et al., 2020). It is also mentioned that for practical cases, computing permutation matrix for different values of $\eta$ in an iterative way can result in better accuracy. The GRAMPA uses $\eta=0.2$ for all their experiments (Fan et al., 2020).
We used a few values of $\eta$ in the classification problem using the MUTAG dataset and got that the best accuracy is obtained at $\eta=0.5$. In Figure 6, the performance of the tuning parameter is demonstrated.
@@ A.9 Effect of tuning parameter $\eta$ on Graph Matching
A.10 Experimental Setup
@@
To evaluate the performance of the Graph Geodesic Distance (GGD) metric, we utilized graph datasets from the TUDataset collection (Morris et al., 2020). For small graphs, we used datasets like MUTAG and BZR, and for larger graphs, we selected PC-3H and SW-620H, which present more sizable networks. Detailed information about the datasets used is provided in Table 10.

### v3

@@ A.9 Choosing $\epsilon$ for converting Laplacians to SPD matrices
A.10 Partial Node Features
Cutting-edge graph distance metrics like TMD rely on node attributes to compute the dissimilarity between graphs, resulting in more accurate outcomes when all attributes are available. However, acquiring datasets with complete node attributes is often unattainable in real-world scenarios, leading to partially missing features In such scenarios when only partial node features are available, we compare TMD with GGD to better understand their differences. Table 6 shows that the TMD metric outperforms GGD at various levels when node features are fully accessible. However, when node features are randomly removed from the MUTAG dataset, the accuracy of TMD degrades substantially.
Dist Metric 0% 20% 50% 80% 100%
GGD 0.78 0.78 0.77 0.77 0.77
TMD, L = 3 0.84 0.78 0.72 0.63 0.61
TMD, L = 4 0.81 0.77 0.62 0.58 0.57
TMD, L = 5 0.80 0.75 0.65 0.58 0.53
A.11 Approximate GGD on Small Graphs Using Extreme Eigenvalues
Table 7: Performance of GGD using extreme eigenvalues only.
Task Number of extreme eigenvalues
2 4 All
Correlation with GNN 0.74 0.76 0.77
[…]
The largest and the smallest eigenvalues correspond to the most dominant mismatches in graph cuts and effective resistance distances, contributing the most to the total GGD value. Similarly, the second largest and smallest eigenvalues correspond to the next significant mismatched cuts. In our experiment, we obtain approximate GGDs using a few extreme eigenvalue pairs and compare them with the ground truth. Figure 7 illustrates the relative accuracy of the approximate GGDs, in which we observe that the top four pairs of extreme eigenvalues contribute $80\%$ of the total GGD values. In addition, we conduct the SVC classification task and GNN correlation study using GGD with only 22 and $4$ extreme eigenvalue pairs, respectively, and present the associated findings in Table 7.
A.12 Scalability of GGD Approximation on Very Large Graphs
For very large graphs, computing the full spectrum of eigenvalues becomes computationally expensive. To address this, we use approximate GGD strategy using only a small fraction of the extreme eigenvalues. Our results in Table 8 show that even with just 2% of the extreme eigenvalues, the approximation remains highly correlated with the original GGD value. Moreover, the percentage required decreases with increasing graph size, significantly improving scalability. Table 9 demonstrates the resulting runtime improvements.
Table 8: Correlation between GGD values and approximate GGD values using a portion of extreme eigenvalues for very large graphs, where n represents the number of nodes.
Extreme Eigenvalues Used 0.5% 1% 2% 4% All
Correlation (n $\in$ [5000, 5200]) 0.8323 0.9361 0.9772 0.9902 1
Correlation (n $\in$ [10000, 10500]) 0.8666 0.9501 0.9808 0.9928 1
Correlation (n $\in$ [15000, 16000]) 0.8892 0.9599 0.9879 0.9952 1
Table 9: Runtime Comparison of Exact vs. Approximate GGD on Very Large Graphs
Node
Numbers Calculation Time
GGD Aprx GGD (2% Eigs) Aprx GGD (4% Eigs)
n $\in$ [5000, 5200] 29.13s 7.11s 7.98s
n $\in$ [10000, 10500] 241.33s 55.56s 56.29s
n $\in$ [15000, 16000] 920.77s 162.30s 163.64s
A.13 Comparison of Two Different Riemannian Metrics for SPD Matrices
The two most commonly used Riemannian metrics on the SPD manifold are the Affine Invariant Riemannian Metric (AIRM) and the Log-Euclidean Riemannian Metric (LERM) [39, 40, 41]. AIRM is a Riemannian metric that remains invariant under affine transformations, meaning the metric is unaffected when matrices are transformed by any invertible operation. The geodesic distance between two SPD matrices, A and B, using AIRM is given by [63, 68]:
@@ A.13 Comparison of Two Different Riemannian Metrics for SPD Matrices
On the other hand, LERM addresses some of the computational complexity challenges associated with AIRM by mapping SPD matrices to an Euclidean space through the matrix logarithm operation. In this Euclidean space, computations are simplified. The geodesic distance between two SPD matrices, A and B, using LERM is defined as [69]:
@@
A.14 Using Normalized Laplacians for GGD Calculation
In many existing studies, the normalized Laplacian matrix is commonly used to study spectral graph properties [70]. The normalized Laplacian matrix of a graph $G$ is defined as: $L_{norm}=I-A_{norm}$, where $A_{norm}$ is the normalized adjacency matrix. The normalized adjacency matrix is expressed as [70]:
@@ A.14 Using Normalized Laplacians for GGD Calculation
GGD using normalized Laplacian matrices 384.097 254.440 188.345 165.332
A.15 Effect of tuning parameter $\eta$ on Graph Matching
In the original work, it was suggested that the regularization parameter $\eta$ needs to be chosen so that $\sigma\vee n^{-0.1}\lesssim\eta\lesssim 1/\log n$ [33]. It is also mentioned that for practical cases, computing permutation matrix for different values of $\eta$ in an iterative way can result in better accuracy. The GRAMPA uses $\eta=0.2$ for all their experiments [33]. We used a few values of $\eta$ in the classification problem using the MUTAG dataset and got that the best accuracy is obtained at $\eta=0.5$. In Figure 9, the performance of the tuning parameter is demonstrated.
@@ A.15 Effect of tuning parameter $\eta$ on Graph Matching
A.16 Sensitivity of GGD to Simple Graph Perturbations
Due to the simplicity and general applicability of the GGD, it can be readily computed between any pair of undirected graphs. To evaluate the robustness of GGD under simple structural changes, we conducted a series of experiments involving basic graph perturbations across both small and large graphs.
Our results show that node and edge additions typically have a minimal impact on the GGD values, suggesting that the metric is largely invariant to the inclusion of redundant elements. In contrast, node and edge drops exhibit noticeable effect. While a significant portion of these perturbations still yield GGD values comparable to the original graphs, some cases show some deviations. These deviations are associated with the removal of structurally important nodes or edges, which alters the graph topology. The quantitative results of these perturbation experiments are summarized Table 12, 13 and Figure 10, 11, 12, 13, 14, 15, 16, 17.
Table 12: Correlation between GGD values before and after applying different perturbation methods on smaller graphs (20–50 nodes)
Amount of Perturbation 1 2 3 4 5
Nodes Dropped 0.6934 0.6084 0.5896 0.5614 0.5368
Nodes Added 0.9976 0.9959 0.9943 0.9922 0.9900
Edges Removed 0.8375 0.7890 0.6644 0.6537 0.6332
Edges Added 0.9990 0.9982 0.9974 0.9964 0.9954
Table 13: Correlation between GGD values before and after applying different perturbation methods on larger graphs (200–500 nodes)
Amount of Perturbation 5 10 15 20 25 30
Nodes Dropped 0.8255 0.7734 0.6533 0.6116 0.5736 0.5563
Nodes Added 0.9991 0.9981 0.9963 0.9944 0.9915 0.9884
Edges Removed 0.9137 0.8857 0.8265 0.7989 0.7550 0.7226
Edges Added 0.9996 0.9992 0.9986 0.9979 0.9971 0.9963
Figure 10: GGD (x-axis) vs GGD after Node Drop perturbation (y-axis) for smaller graphs.
Figure 11: GGD (x-axis) vs GGD after Node Drop perturbation (y-axis) for larger graphs.
Figure 12: GGD (x-axis) vs GGD after Node Addition perturbation (y-axis) for smaller graphs.
Figure 13: GGD (x-axis) vs GGD after Node Addition perturbation (y-axis) for larger graphs.
Figure 14: GGD (x-axis) vs GGD after Edge Drop perturbation (y-axis) for smaller graphs.
Figure 15: GGD (x-axis) vs GGD after Edge Drop perturbation (y-axis) for larger graphs.
Figure 16: GGD (x-axis) vs GGD after Edge Addition perturbation (y-axis) for smaller graphs.
Figure 17: GGD (x-axis) vs GGD after Edge Addition perturbation (y-axis) for larger graphs.
A.17 Graph formation from Dataset using Probabilistic Graphical Models
We take similiar type of datasets such as MNIST, Fashion-MNIST, KMNIST, and USPS [58] and convert them into connected graphs. To construct these graph structures, we use Probabilistic Graphical Models (PGMs), also known as Markov Random Fields (MRFs) [71]. PGMs are powerful tools in machine learning and statistical physics for representing complex systems with intricate dependency structures [72]. They encode the conditional dependencies between random variables through an undirected graph structure. Recent studies have shown that the graph structure learned via PGMs can exhibit resistance distances that encode the Euclidean distances between their corresponding data samples [73].
We create a Feature Matrix (FM) $U$ from the dataset, where each row represents a data sample, and the row $U_{p}$ itself serves as the feature vector of that sample $p$, on this context- the pixel values. A dense $k$-nearest neighbor (k-NN) graph $G_{dense}$ is initially constructed using the FM. To obtain the final dataset graph $G_{dataset}$, spectral sparsification is applied by solving the convex optimization problem [71]:
$\max_{\Theta}\quad F(\Theta)=\log\operatorname{det}(\Theta)-\frac{1}{k}\operatorname{Tr}\left(U^{\top}\Theta U\right)$ (29)
where $\Theta=L+\frac{1}{\sigma^{2}}I$. Here, $L$ is the graph Laplacian, $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix, $I$ is the identity matrix, and $\sigma^{2}>0$ represents the prior feature variance. To solve this the following lemma is used:
Lemma A.3.
Maximizing the objective function in Equation 29 can be achieved in nearly-linear time via the following edge pruning strategy equivalent to spectral sparsification of the initial dense nearest-neighbor graph. Specifically, edges with small distance ratios
$\rho_{p,q}=\frac{R_{\mathrm{eff}}(p,q)}{d_{\mathrm{euc}}(p,q)}=w_{p,q}\cdot R_{\mathrm{eff}}(p,q)$ (30)
are pruned, where $R_{\mathrm{eff}}(p,q)$ denotes the effective resistance distance between nodes $p$ and $q$, $d_{\mathrm{euc}}(p,q)=\left\|U_{p}-U_{q}\right\|_{2}^{2}$ represents the data distance between the feature of nodes $p$ and $q$, and $w_{p,q}=\frac{1}{d_{\mathrm{euc}}(p,q)}$ is the weight of edge $(p,q)$ [71].
Computing the edge sampling probability $\rho_{p,q}$ for each edge $(p,q)$ becomes computationally expensive for large graphs. To address this, an improved algorithm using a low-resistance-diameter (LRD) decomposition is proposed, extending the short-cycle decomposition [74] to weighted graphs. The method efficiently computes effective resistance to partition the graph into clusters, thereby enhancing the sparsification process. This results in a low-dimensional graph $G_{dataset}$ that retains important structural properties while reducing dimensionality [71].
A.18 Experimental Setup
@@
To evaluate the performance of the Graph Geodesic Distance (GGD) metric, we utilized graph datasets from the TUDataset collection [35]. For small graphs, we used datasets like MUTAG and BZR, and for larger graphs, we selected PC-3H and SW-620H, which present more sizable networks. Detailed information about the datasets used is provided in Table 14.